Two gasohol mixtures are available, one with 6.35% alcohol and the other with 11.28% alcohol. How many litres of the 6.35% mixture must be added to 25.0 L of the other mixture to make a final mixture containing 8.00% alcohol?
Round the answer to one decimal place

Answer by Boreal(15235)

x=number of l needed.
.0635x+(25)(.1128)=(25+x)(.08), the "pure" amount
.0635+2.82=2+.08x
-0.0165x=-.82
x=49.7 liters.
Rough check.
This is about twice as much of the lower amount, and the lower amount is 1.65% less concentrated and the higher amount is 3.3% more concentrated. This makes sense.